OpenStudy (anonymous):

Log₂(3*2^(x-1)-1)=2x-1 Can't seem to reach the answer (1 and 0)

OpenStudy (anonymous):

\[\log_{2} (3\times2^{x-1}-1)=2x-1 \] What I tried so far: 3*2^(x-1)-1=2^(2x-1) 3*2^(x)*2^(-1)-1=2^(2x)*2^(-1) 2^(x) = t 3*t*2^(-1)-1=t^(2)*2^(-1) t*(1/2)=t^(2)*(1/2) /: 1/2 , /: t t=0, t=1 2^x=0, 2^x=1 no solution, 2^x=2^0 x=0

OpenStudy (anonymous):

i hope this helps http://www.mathportal.org/calculators.php

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ \log_2(3\times 2^{x-1}-1)=2x-1\LARGE\color{white}{ \rm │ }}\)

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ \log_2(3\times 2^{x-1}-1)=2x-1\LARGE\color{white}{ \rm │ }}\) \(\normalsize\color{blue}{ \log_2(3\times 2^{x-1}-1)=(2x-1)\log_2(2)\LARGE\color{white}{ \rm │ }}\) \(\normalsize\color{blue}{ \log_2(3\times 2^{x-1}-1)=\log_2(2^{2x-1})\LARGE\color{white}{ \rm │ }}\) \(\normalsize\color{blue}{ 3\times 2^{x-1}-1=2^{2x-1}\LARGE\color{white}{ \rm │ }}\)

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ (3/2)2^{x}-1=(1/2)2^{x+x} \LARGE\color{white}{ \rm │ }}\) \(\normalsize\color{blue}{ (3)2^{x}-2=2^{x+x} \LARGE\color{white}{ \rm │ }}\) is what I am up to so far.

OpenStudy (solomonzelman):

Pilot, please share what you got, because I am not very sure...

OpenStudy (anonymous):

let u = 2^x and solve as a quadratic

OpenStudy (anonymous):

3u-2=2u^2 => 2u^2 - 3u +2 = 0 => (2u-1)(u-1) => u = 1 and u = 1/2 so 2^x = 1 and 2^x = 1/2 => x = 0 and x = -1

OpenStudy (anonymous):

but -1 can't be an answer because log can't = -. so either i goofed on factoring or solutions given are incorrect.

OpenStudy (anonymous):

I tried inputting it into wolfram, they seem to be in order http://www.wolframalpha.com/input/?i=Log2%283*2%5E%28x-1%29-1%29%3D2x-1

OpenStudy (anonymous):

but i just can't tell where's the mistake

OpenStudy (anonymous):

0 can't be a solution because it leads to an false statement. hold on 1 sec...

OpenStudy (anonymous):

x=0 in the original equation: log2(3*2^(0-1)-1) = 2*0-1 log2(1/2) = -1 -1 = -1

OpenStudy (anonymous):

\[3\cdot 2^{x-1}-1=2^{2x-1} \Rightarrow \frac{3}{2}2^x-1=\frac{1}{2}2^{2x} \Rightarrow \frac{1}{2}\left(2^x \right)^2-\frac{3}{2}\left( 2^x \right)-1\]Let \(u=2^{x}\)\[\Rightarrow u^2-3u+2 = 0 \Rightarrow \left( u-2 \right)\left( u-1 \right)=0 \Rightarrow u=2 \text{ , } u=1\]\[\Rightarrow 2^x=2 \text{ and }2^x=1\Rightarrow x=1 \text{ and } x=0\] But \(x\ne 0\) because RHS is then negative in the original equation. So x=1 is the only solution.

OpenStudy (anonymous):

oops, disregard the last part about the negative. the argument of log can't be negative, not the value.

OpenStudy (anonymous):

logs are just exponents and negative exponents are perfectly okay!

OpenStudy (solomonzelman):

Nice idea, but for u, 3x would be ?

OpenStudy (solomonzelman):

It won't be just 3u, will it ?

OpenStudy (anonymous):

3x? it should be 3*2^x

OpenStudy (anonymous):

u=2^x

OpenStudy (solomonzelman):

no, you got it...

OpenStudy (anonymous):

no, the original is \[\log_{2} \left( 3\cdot 2^{x-1}-1 \right)=2x-1\]

OpenStudy (anonymous):

all good? there is a typo... should be \[\frac{1}{2}(2^x)^2−\frac{3}{2}(2^x)+1\] not\[\frac{1}{2}(2^x)^2−\frac{3}{2}(2^x)-1\]

OpenStudy (anonymous):

yeah, great find pilot. I'm going to use your way to solve the other similar exercises. much appriciated.

OpenStudy (anonymous):

you're welcome! glad i could help!

OpenStudy (anonymous):

it's a fairly common "trick" but not always obvious. in calc, it's used all the time but kinda in reverse; anytime you do the chain rule with derivatives.